From a box containing 6 black balls and 4 green balls, 3 balls are drawn in succession, each ball being replaced in the box before the next draw is made.
a) all 3 are the same color?
b) each color is represented?

Answer by stanbon(57361)

From a box containing 6 black balls and 4 green balls, 3 balls are drawn in succession, each ball being replaced in the box before the next draw is made.
a) all 3 are the same color?
Pick 3 black: 6C3 = 20 ways
Pick 3 green: 4C3 = 4 ways
----
Pick 3 randomly = 10C3 = 120
P(3 of same color) = 24/120
-------------------------------------
b) each color is represented?
P(each color is represented)
= 1- P(3 of same color)
= 1-(24/120)
= 96/120
==============
Cheers,
Stan H.
==============